OpenStudy (anonymous):

also I am stuck on 4^x+6=3^3-4 I have to round to the nearest app. decimal I got the equation in but am not getting it maybe put in my calculator right idk

OpenStudy (anonymous):

it is exponential and logarithmic functions

OpenStudy (anonymous):

4^x = 27 -4-6 4^x =17 ln(4^x) = ln (17) x*ln(4)= ln(17) x=ln(17)/ln(4) x=2.04 does it help?

OpenStudy (anonymous):

that is the wrong answer they had-44.1884